It is known that a quadratic function

f(x)=ax^2+bx+2 (a<0)

satisfies the relation f(10)=0. Find the largest possible quantity of integer solutions of the inequality ax^4+bx^2+2>0.


1
Expert's answer
2021-09-24T11:03:31-0400

"f(x) = a x\\^2 + bx + 2 \ (a<0)". "a \ is \ negative." Therefore, the function can be written as: "f(x) = -a x\\^2 + bx + 2"

"f(10) = -a(10)\\^2 + b(10) + 2= 0"

"-100a + 10b + 2 = 0"

"10 b= 100a - 2"

"b= 10 a- \frac{2}{10}"

"ax\\^4 + bx\\^2 + 2>0"

"ax\\^4 + (10a - \frac{2}{10}) x\\^2 + 2 >0"

"f(10) = 0"

"a10\\^4 + (10a - \frac{2}{10}) 10\\^2 + 2 >0"

"10000a + (10a - \frac{2}{10}) 100 + 2 >0"

"10000a + 1000a -20 + 2 >0"

"11000a > 18"

"a>\frac{18}{11000}"

From the above the, largest possible quantity of integer is